# Adaptive Beam Alignment using Noisy Twenty Questions Estimation with Trained Questioner

### v1

@@ Abstract
The 6G communication systems use millimeter-wave and multiple-input multiple-output technologies to achieve wide bandwidth and high throughout, leading to indispensable need for beam alignment to overcome severe signal attenuation. Traditional sector-search-based beam alignment algorithms rely on sequential sampling to identify the best sector, resulting in a significant latency burden on 6G communication systems. Recently proposed adaptive beam alignment algorithms based on the active learning framework address the problem, aiming to identify the optimal sector with the fewest possible samples under an identical sector partition. Nevertheless, these algorithms either lack feasibility (Chiu, Ronquillo and Javidi, JSAC 2019) due to ideal assumptions or lack interpretability (Sohrabi, Chen and Yu, JSAC 2021) due to the use of end-to-end black-box neural networks. To avoid ideal assumptions and maintain interpretability, in this paper, we address all above problems by proposing an adaptive beam alignment algorithm using the framework of noisy twenty questions estimation with a trained questioner. Specifically, we use two methods for training the questioner to eliminate reliance on ideal assumptions. The first method maps queries of twenty questions estimation to beamforming vectors via weighted summation of steering vectors, as an initial attempt to address the feasibility problem encountered in prior pioneering study by Chiu, Ronquillo and Javidi (JSAC 2019). The second method uses multi-layer fully connected neural networks to achieve improved performance while only employing them to train the questioner, which can effectively mitigate the interpretability issues in prior study by Sohrabi, Chen and Yu (JSAC 2021). Furthermore, we provide numerical simulations to illustrate the effectiveness of our proposed adaptive beam alignment algorithms and demonstrate that our algorithms outperform all benchmark algorithms.
@@ I Introduction
Traditional sector-search-based BA algorithms include naive exhaustive search and hierarchical exhaustive search with a fixed codebook of given beam directions. In the naive exhaustive BA algorithm, the transmitter (TX) periodically sends pilot signals, and the receiver (RX) selects the optimal beam direction based on the received pilot signals using criteria such as the maximum received signal strength indicator. However, this algorithm needs to sample each sector sequentially, which results in a significant latency burden. Hierarchical exhaustive BA algorithms, such as those specified by IEEE 802.11 ad [10], IEEE 802.15 3c [9] and 5G NR [37], partially address the above problem by progressively reducing, over multiple stages, the number of sectors that need to be searched. Nevertheless, since the codebook is fixed rather than adaptively generated, the hierarchical approach provides only limited performance improvement and suffers from error propagation.
@@
In order to minimize the number of samples, Chiu, Ronquillo and Javidi [7] adopted the idea of active learning, specifically noisy twenty questions estimation pioneered by Rényi [22] and Ulam [29]. Twenty questions estimation is a question and answer style game between two players: one player called the oracle picks something as a secret and the other player called the questioner aims to accurately guess the secret by posing as fewer queries as possible. In the BA algorithm of [7], at each time point, designing a beamforming vector corresponds to designing a query in twenty questions estimation while one measurement result of the received signal corresponds to a response from an oracle in twenty questions. Two measurement rules are considered in [7]: full measurement and 1-bit measurement. With full measurement, one uses the complete information of received signals to estimate the angle-of-arrival (AoA), while with 1-bit measurement, only extreme quantized information of received signals can be obtain. Using the connection between BA and twenty questions estimation, Chiu, Ronquillo and Javidi et al. used a hierarchical posterior matching strategy to generate queries, converted queries into beamforming vectors with a pre-designed ideal beamforming codebook, and estimated the AoA using noisy responses and decoding function of twenty questions. Subsequently, many studies applied the idea of active learning to BA via other tools of machine learning, such as neural network (NN) [24, 12, 4] and multi-armed bandits (MAB) [31, 32, 15, 2, 3]. In particular, the authors integrated the main steps of BA, including the design of beamforming vectors, the update of posterior probability vectors and the estimation of AoA, into a deep neural network (DNN) for training. The authors of [12] focused on the design of beamforming vector using a recurrent neural network (RNN) with long short-term memory. In [31], the authors proposed a hierarchical exhaustive algorithm using multi-armed bandits by modeling each narrow beam direction as a base arm, combining several adjacent base arms as a super arm with wide beam direction and proposing a two-phase track-and-stop algorithm that firstly pulls super arms and then pulls base arms.
@@
However, all above algorithms have limitations. Specifically, the BA algorithm in [7] did not specify how the ideal beamforming vector in [7, Assumption 4] is designed, making it difficult to be applied in practical communication systems. While the studies in [24, 12] addressed the above feasibility problem, the extensive use of DNNs or RNNs reduces the interpretability of the algorithms, in the sense that their performance can neither be explained by existing theoretical results in [6] nor supported by new theoretical analyses. Although the algorithms based on MAB [31, 32] had good interpretability, it requires frequent arm pulling, which implies frequent sector searches, thereby leading to undesired high latency overhead. In summary, existing algorithms suffered the problem of feasibility, interpretability or high latency. One might wonder whether it is possible to propose a feasible and interpretable BA algorithm that has low latency and good performance. In this paper, we answer the question affirmatively. Our main contributions are summarized as follows.
@@ I-A Main Contributions
  • Inspired by [24], we employ a DNN to facilitate the transformation from queries to beamforming vectors, with the goal of achieving less performance degradation compared with the LWS-based method. Specifically, we calculate the steering vectors of angles within the search region and feed them as inputs into a multi-layer fully connected NN. Compared with the LWS-based method, the DNN-based method achieves better performance in numerical simulations. Furthermore, compared with BA algorithm in [24] that uses the end-to-end DNN, our algorithm improves interpretability by limiting the role of the DNN in the active-learning-based BA algorithm.
  • We next investigate the following questions via numerical simulations in Figs. 9-14 and analyses of the two proposed methods: what limits active-learning-based beam alignment algorithms from achieving their theoretical optimum; why the trained questioner based on DNN method and the full measurement rule can exhibit better performance; and whether the size of the search region affects the performance of the trained questioner.
  • Finally, we run comprehensive numerical examples to demonstrate the performance of our BA algorithms in Fig. 15. In our numerical simulations, we first show that the performance of our BA algorithm under the 1-bit measurement rule is better than the the naive beam sweeping algorithm and hierarchical beam sweeping algorithm [37] under the full measurement rule. Furthermore, under the full measurement rule, we run extensive numerical examples to compare the performance our algorithms, the naive beam sweeping algorithm, hierarchical beam sweeping algorithm [37] and the end-to-end DNN-based BA algorithm [24]. Simulation results show that i) the performance of our BA algorithms under the full measurement rule is much better than under the 1-bit measurement rule and ii) under the full measurement rule, our proposed BA algorithms outperform all benchmarking BA algorithms.
@@ II-A Adaptive Beam Alignment
where $\mathbf{n}_{t}\sim\mathcal{CN}(\{\mathbf{0}\}_{N_{R}},\sigma^{2}\mathbf{I}_{N_{R}})$ is the independent Gaussian noise vector with homogeneous variance $\sigma^{2}\in\mathbb{R}_{+}$. To facilitate the subsequent analysis, we temporarily assume $\alpha=P=1$. In Section IV-D, we will address the case of unknown $\alpha$ using a Kalman filter.
@@
The full measurement rule leads to better performance, as the received signal can be fully used, and serves as a representative of ideal communication conditions; while the 1-bit measurement rule simplifies the theoretical analysis, and represents one of the most extreme communication conditions.
@@ II-B Noisy Adaptive Twenty Questions Estimation
This subsection introduces the noisy twenty questions estimation framework [33]. Let $\mathbf{S}=(S_{1},\ldots,S_{l})\in[0,1]^{l}$ be a $l$-dimensional random variable generated from an arbitrary distribution $f_{\mathbf{S}}\in\mathcal{F}(\mathcal{S})$. As shown in Fig. 4, a questioner aims to estimate a random variable $\mathbf{S}$ with the assistance of an oracle who has access to $\mathbf{S}$. At each time point $t\in\mathbb{N}$, the questioner generates a Lebesgue measurable query $\mathcal{A}_{t}\subseteq[0,1]^{l}$ and asks the oracle whether $\mathbf{S}$ lies in the set $\mathcal{A}_{t}$. The query set $\mathcal{A}_{t}$ is generated using all previous queries $(\mathcal{A}_{1},\ldots,\mathcal{A}_{t-1})$ and their noisy responses $(Y_{1},\ldots,Y_{t-1})$. After receiving $\mathcal{A}_{t}$, the oracle generates binary yes/no answers $X_{t}=\mathds{1}(\mathbf{S}\in\mathcal{A}_{t})$ and passes $X_{t}$ over a query-dependent channel [34, Eq. (1)] with transition probability $P_{Y_{t}|X_{t}}^{\beta(|\mathcal{A}_{t}|)}\in\mathcal{P}(\mathcal{Y}|\mathcal{X})$, yielding a noisy response $Y_{t}$, where $\beta(\cdot)$ is a bounded Lipschitz continuous function, and $|\mathcal{A}_{t}|$ denotes the size of the query $\mathcal{A}_{t}$. For example, a query-dependent binary symmetric channel (BSC) is defined as follows.
@@ III-A Linear Weighted Sum
With the available mapping function from search regions to beamforming vectors proposed in Definition 3, now we can provide a feasible adaptive BA algorithm, summarized in Algorithm 1. To evaluate the performance of the methods of LWS, we design numerical simulations with 1-bit measurement rule. At each time point $t$, we calculate the power of each received signal coming from the angle corresponding to the midpoint of the secondary sub-interval within the search region $\mathcal{A}_{t}$. Unless otherwise specified, in all subsequent numerical simulations, we select the minimum $|\mathbf{y}_{\mathrm{min}}^{\mathrm{q}}|^{2}$ as the power threshold $\iota_{t}$. Specifically, we define the derived variable $K:=|\mathcal{A}_{t}|M$, which denotes the number of sub-intervals within the search region $\mathcal{A}_{t}$. For any search region $\mathcal{A}_{t}\subseteq[0,1]$ and $k^{\prime}\in[K\times k]$, let $\theta_{k^{\prime}}$ denote the angle corresponding to the midpoint of the $k^{\prime}$-th secondary sub-interval within the search region, and let $\mathbf{y}_{\theta_{k^{\prime}}}^{\mathrm{q}}$ denote the received signal from the angle $\theta_{k^{\prime}}$. The definition of $|\mathbf{y}_{\mathrm{min}}^{\mathrm{q}}|^{2}$ is provided as follows:
@@
Unless otherwise specified, this way of threshold selection will be consistently adopted. We calculate the power of each received signal coming from the angle corresponding to the midpoint of the secondary sub-interval within the non-search region and record the maximum $|\mathbf{y}_{\mathrm{max}}^{\mathrm{nonq}}|^{2}$, whose definition is similar to $|\mathbf{y}_{\mathrm{min}}^{\mathrm{q}}|^{2}$ in (16) as follows:
@@ IV-B3 Ideal hiePM-based strategy proposed in [7]
where $\mathcal{D}_{l}^{k}$ denotes the probing region, $C_{s}$ denotes a normalization constant, $\mathbf{A}_{\text{BS}}$ denotes a matrix of array manifolds, and $\mathbf{G}_{\mathcal{D}_{l}^{k}}$ denotes a vector indicating the probing region. The term $\sigma_{0}^{2}\mathbf{I}$ we added here is for diagonal loading, where $\sigma_{0}^{2}$ denotes a normalization factor, and $\mathbf{I}$ denotes the identity matrix. Furthermore, the threshold selection of this method follows the approach described in [7, Appendix A].
@@ IV-C Simulation Results
(a) Average quadratic loss versus SNR for the two proposed methods with 1-bit measurement rule, the ideal hiePM-based method, and the two beam sweeping algorithms.
@@ IV-D The Case of Unknown Fading Coefficient $\alpha$
Let $\mathbf{c}_{i,t}=\mathbf{w}_{t}^{H}\mathbf{A}(\theta_{i})$. Assuming that the unknown fading coefficient $\alpha$ follows a complex Gaussian distribution with mean $\mu_{\alpha}$ and variance $\sigma_{\alpha}$. Recall the definition of the received signal $\mathbf{y}_{t}$ in (3) at any time point $t\in\mathbb{N}$. For any presumed AoA $\theta_{i}\in[\theta_{\text{min}},\theta_{\text{max}}]$, the corresponding mean $\mu_{i,t}^{\alpha}$ and variance $\sigma_{i,t}^{\alpha}$ of $\alpha$ can be updated by Kalman filter [14]:
@@
To demonstrate the feasibility of the Kalman filter estimator on the two proposed methods, we also plot the average quadratic loss $\mathbb{E}[(\hat{\theta}-\theta)^{2}]$ with full measurement rule under the setting of $\alpha\sim\mathcal{CN}(1,0.25^{2})$ over $10^{3}$ Monte Carlo runs. For the consistency, we only replace the definition of the probability factor $\nu$ in (20) with (24). As shown in Fig. 16, both two estimator are applicable to with acceptable performance loss. Furthermore, as a measure of storage and computational complexity, Fig. 17 shows that the average runtime of the Kalman filter estimator is significantly less than the MMSE estimator for the same average quadratic loss, which is consistent with the analysis of the advantages of Kalman filter estimator in [24].

### v2

@@ Abstract
The 6G communication systems use millimeter-wave and multiple-input multiple-output technologies to achieve wide bandwidth and high throughput, leading to an indispensable need for beam alignment to overcome severe signal attenuation. Traditional sector-search-based beam alignment algorithms rely on sequential sampling to identify the best sector, resulting in a significant latency burden on 6G communication systems. Recently proposed adaptive beam alignment algorithms based on the active learning framework address the problem, aiming to identify the optimal sector with the fewest possible samples under the same sector partition. Nevertheless, these algorithms either lack feasibility (Chiu, Ronquillo and Javidi, JSAC 2019) due to ideal assumptions or lack interpretability (Sohrabi, Chen and Yu, JSAC 2021) due to the use of end-to-end black-box neural networks. To avoid ideal assumptions and maintain interpretability, in this paper, we address all above problems by proposing an adaptive beam alignment algorithm using the framework of noisy twenty questions estimation with a trained questioner. Specifically, we use two methods for training the questioner to eliminate reliance on ideal assumptions. The first method maps queries of twenty questions estimation to beamforming vectors via weighted summation of steering vectors, as an initial attempt to address the feasibility problem encountered in prior pioneering study by Chiu, Ronquillo and Javidi (JSAC 2019). The second method uses multi-layer fully connected neural networks to achieve improved performance while only employing them to train the questioner, which can effectively mitigate the interpretability issues in prior study by Sohrabi, Chen and Yu (JSAC 2021). Furthermore, we provide numerical simulations to illustrate the effectiveness of our proposed adaptive beam alignment algorithms and demonstrate that our algorithms outperform all benchmark algorithms.
@@ I Introduction
Traditional sector-search-based BA algorithms include naive exhaustive search and hierarchical exhaustive search with a fixed codebook of given beam directions. In the naive exhaustive BA algorithm, the transmitter (TX) periodically sends pilot signals, and the receiver (RX) selects the optimal beam direction based on the received pilot signals using criteria such as the maximum received signal strength indicator. However, this algorithm needs to sample each sector sequentially, which results in a significant latency burden. Hierarchical exhaustive BA algorithms, such as those specified by IEEE 802.11 ad [10], IEEE 802.15.3c [9] and 5G NR [37], partially address the above problem by progressively reducing, over multiple stages, the number of sectors that need to be searched. Nevertheless, since the codebook is fixed rather than adaptively generated, the hierarchical approach provides only limited performance improvement and suffers from error propagation.
@@
To minimize the number of samples, Chiu, Ronquillo and Javidi [7] adopted the idea of active learning, specifically noisy twenty questions estimation pioneered by Rényi [22] and Ulam [29]. Twenty questions estimation is a question and answer style game between two players: one player called the oracle picks something as a secret and the other player called the questioner aims to accurately guess the secret by posing as fewer queries as possible. In the BA algorithm of [7], at each time point, designing a beamforming vector corresponds to designing a query in twenty questions estimation and one measurement result of the received signal corresponds to a response from an oracle in twenty questions. Two measurement rules are considered in [7]: full measurement and 1-bit measurement. With full measurement, one uses the complete information of received signals to estimate the angle-of-arrival (AoA), while with 1-bit measurement, only extreme quantized information of received signals can be obtained. Using the connection between BA and twenty questions estimation, Chiu et al. used a hierarchical posterior matching strategy to generate queries, converted queries into beamforming vectors with a pre-designed ideal beamforming codebook, and estimated the AoA using noisy responses and decoding function of twenty questions. Subsequently, many studies applied the idea of active learning to BA via other tools of machine learning, such as neural network (NN) [24, 12, 4] and multi-armed bandits (MAB) [31, 32, 15, 2, 3]. In particular, the authors integrated the main steps of BA, including the design of beamforming vectors, the update of posterior probability vectors and the estimation of AoA, into a deep neural network (DNN) for training. The authors of [12] focused on the design of beamforming vector using a recurrent neural network (RNN) with long short-term memory. In [31], the authors proposed a hierarchical exhaustive algorithm using multi-armed bandits by modeling each narrow beam direction as a base arm, combining several adjacent base arms as a super arm with wide beam direction and proposing a two-phase track-and-stop algorithm that firstly pulls super arms and then pulls base arms.
@@
However, all the above algorithms have limitations. Specifically, the BA algorithm in [7] did not specify how the ideal beamforming vector in [7, Assumption 4] is designed, making it difficult to apply in practical communication systems. Although the studies in [24, 12] addressed the above feasibility problem, the extensive use of DNNs or RNNs reduces the interpretability of the algorithms, in the sense that their performance can neither be explained by existing theoretical results in [6] nor supported by new theoretical analyses. Although the algorithms based on MAB [31, 32] had good interpretability, they require frequent arm pulling, which implies frequent sector searches, thereby leading to undesired high latency overhead. In summary, existing algorithms suffer from issues of feasibility, interpretability or high latency. One might wonder whether it is possible to propose a feasible and interpretable BA algorithm that has low latency and good performance. In this paper, we answer the question affirmatively. Our main contributions are summarized as follows.
@@ I-A Main Contributions
  • Inspired by [24], we employ a DNN to facilitate the transformation from queries to beamforming vectors, with the goal of achieving less performance degradation compared with the LWS-based method. Specifically, we calculate the steering vectors of angles within the search region and feed them as inputs into a multi-layer fully connected NN. Compared with the LWS-based method, the DNN-based method achieves better performance in numerical simulations. Furthermore, compared with the BA algorithm in [24] which adopts an end-to-end DNN architecture, our algorithm improves interpretability by limiting the role of the DNN in the active-learning-based BA algorithm.
  • We next investigate the following questions via numerical simulations of the two proposed methods in Figs. 9-14: what limits active-learning-based beam alignment algorithms from achieving their theoretical optimum; why the questioner trained by the DNN-based method and the full measurement rule can exhibit better performance; and whether the size of the search region affects the performance of the trained questioner.
  • Finally, we run comprehensive numerical examples to demonstrate the performance of our BA algorithms in Fig. 15. In our numerical simulations, we first show that the performance of our BA algorithm under the 1-bit measurement rule is better than the naive beam sweeping algorithm and hierarchical beam sweeping algorithm [37] under the full measurement rule. Furthermore, under the full measurement rule, we run extensive numerical examples to compare the performance of our algorithms, the naive beam sweeping algorithm, hierarchical beam sweeping algorithm [37] and the end-to-end DNN-based BA algorithm [24]. Simulation results show that i) the performance of our BA algorithms under the full measurement rule is much better than under the 1-bit measurement rule and ii) under the full measurement rule, our proposed BA algorithms outperform all benchmark BA algorithms.
@@ II-A Adaptive Beam Alignment
where $\mathbf{n}_{t}\sim\mathcal{CN}(\{\mathbf{0}\}_{N_{R}},\sigma^{2}\mathbf{I}_{N_{R}})$ is the independent Gaussian noise vector with homogeneous variance $\sigma^{2}\in\mathbb{R}_{+}$. To facilitate the subsequent analysis, we temporarily assume $\alpha=P=1$. In Section IV-D, we will address the case of unknown $\alpha$ using a Kalman filter estimator.
@@
The full measurement rule, serving as a representative of ideal communication conditions, leads to better performance as the received signal can be fully used, while the 1-bit measurement rule simplifies the theoretical analysis and represents one of the most extreme communication conditions.
@@ II-B Noisy Adaptive Twenty Questions Estimation
This subsection introduces the noisy twenty questions estimation framework [33]. Let $\mathbf{S}=(S_{1},\ldots,S_{l})\in[0,1]^{l}$ be an $l$-dimensional random variable generated from an arbitrary distribution $f_{\mathbf{S}}\in\mathcal{F}(\mathcal{S})$. As shown in Fig. 4, a questioner aims to estimate a random variable $\mathbf{S}$ with the assistance of an oracle who has access to $\mathbf{S}$. At each time point $t\in\mathbb{N}$, the questioner generates a Lebesgue measurable query $\mathcal{A}_{t}\subseteq[0,1]^{l}$ and asks the oracle whether $\mathbf{S}$ lies in the set $\mathcal{A}_{t}$. The query set $\mathcal{A}_{t}$ is generated using all previous queries $(\mathcal{A}_{1},\ldots,\mathcal{A}_{t-1})$ and their noisy responses $(Y_{1},\ldots,Y_{t-1})$. After receiving $\mathcal{A}_{t}$, the oracle generates binary yes/no answers $X_{t}=\mathds{1}(\mathbf{S}\in\mathcal{A}_{t})$ and passes $X_{t}$ over a query-dependent channel [34, Eq. (1)] with transition probability $P_{Y_{t}|X_{t}}^{\beta(|\mathcal{A}_{t}|)}\in\mathcal{P}(\mathcal{Y}|\mathcal{X})$, yielding a noisy response $Y_{t}$, where $\beta(\cdot)$ is a bounded Lipschitz continuous function, and $|\mathcal{A}_{t}|$ denotes the size of the query $\mathcal{A}_{t}$. For example, a query-dependent binary symmetric channel (BSC) is defined as follows.
@@ III-A Linear Weighted Sum
With the available mapping function from search regions to beamforming vectors proposed in Definition 3, now we can provide a feasible adaptive BA algorithm, summarized in Algorithm 1. To evaluate the performance of the methods of LWS, we design numerical simulations with 1-bit measurement rule. At each time point $t$, we calculate the power of each received signal coming from the angle corresponding to the midpoint of the secondary sub-interval within the search region $\mathcal{A}_{t}$. Specifically, we define the derived variable $K:=|\mathcal{A}_{t}|M$, which denotes the number of sub-intervals within the search region $\mathcal{A}_{t}$. For any search region $\mathcal{A}_{t}\subseteq[0,1]$ and $k^{\prime}\in[K\times k]$, let $\theta_{k^{\prime}}$ denote the angle corresponding to the midpoint of the $k^{\prime}$-th secondary sub-interval within the search region, and let $\mathbf{y}_{\theta_{k^{\prime}}}^{\mathrm{q}}$ denote the received signal from the angle $\theta_{k^{\prime}}$. The definition of $|\mathbf{y}_{\mathrm{min}}^{\mathrm{q}}|^{2}$ is provided as follows:
@@
Unless otherwise specified, in all subsequent numerical simulations, we select the minimum $|\mathbf{y}_{\mathrm{min}}^{\mathrm{q}}|^{2}$ as the power threshold $\iota_{t}$. We calculate the power of each received signal coming from the angle corresponding to the midpoint of the secondary sub-interval within the non-search region and record the maximum $|\mathbf{y}_{\mathrm{max}}^{\mathrm{nonq}}|^{2}$, whose definition is similar to $|\mathbf{y}_{\mathrm{min}}^{\mathrm{q}}|^{2}$ in (16) as follows:
@@ IV-B3 Ideal hiePM-based strategy proposed in [7]
where $\mathcal{D}_{l}^{k}$ denotes the probing region, $C_{s}$ denotes a normalization constant, $\mathbf{A}_{\text{BS}}$ denotes a matrix of array manifolds, and $\mathbf{G}_{\mathcal{D}_{l}^{k}}$ denotes a vector indicating the probing region. The term $\sigma_{0}^{2}\mathbf{I}$ we added here is for diagonal loading, where $\sigma_{0}^{2}$ denotes a normalization factor and $\mathbf{I}$ denotes the identity matrix. Furthermore, the threshold selection of this strategy follows the approach described in [7, Appendix A].
@@ IV-C Simulation Results
(a) Average quadratic loss versus SNR for the two proposed methods with 1-bit measurement rule, the ideal hiePM-based strategy, and the two beam sweeping algorithms.
@@ IV-D The Case of Unknown Fading Coefficient $\alpha$
Let $\mathbf{c}_{i,t}=\mathbf{w}_{t}^{H}\mathbf{A}(\theta_{i})$. Assume that the unknown fading coefficient $\alpha$ follows a complex Gaussian distribution with mean $\mu_{\alpha}$ and variance $\sigma_{\alpha}$. Recall the definition of the received signal $\mathbf{y}_{t}$ in (3) at any time point $t\in\mathbb{N}$. For any presumed AoA $\theta_{i}\in[\theta_{\text{min}},\theta_{\text{max}}]$, the corresponding mean $\mu_{i,t}^{\alpha}$ and variance $\sigma_{i,t}^{\alpha}$ of $\alpha$ can be updated by Kalman filter [14]:
@@
To demonstrate the feasibility of the Kalman filter estimator on the two proposed methods, we also plot the average quadratic loss $\mathbb{E}[(\hat{\theta}-\theta)^{2}]$ with full measurement rule under the setting of $\alpha\sim\mathcal{CN}(1,0.25^{2})$ over $10^{3}$ Monte Carlo runs. For the consistency, we only replace the definition of the probability factor $\nu$ in (20) with (24). As shown in Fig. 16, both estimator are applicable to with acceptable performance loss. Furthermore, as a measure of storage and computational complexity, Fig. 17 shows that the average runtime of the Kalman filter estimator is significantly less than the MMSE estimator for the same average quadratic loss, which is consistent with the analysis of the advantages of Kalman filter estimator in [24].